# Deep-Graph-Sprints: Accelerated Representation Learning in Continuous-Time Dynamic Graphs

### v1

@@ Abstract
Continuous-time dynamic graphs (CTDGs) are essential for modeling interconnected, evolving systems. Traditional methods for extracting knowledge from these graphs often depend on feature engineering or deep learning. Feature engineering is limited by the manual and time-intensive nature of crafting features, while deep learning approaches suffer from high inference latency, making them impractical for real-time applications. This paper introduces Deep-Graph-Sprints (DGS), a novel deep learning architecture designed for efficient representation learning on CTDGs with low-latency inference requirements. We benchmark DGS against state-of-the-art feature engineering and graph neural network methods using five diverse datasets. The results indicate that DGS achieves competitive performance while improving inference speed up to 12x compared to other deep learning approaches on our tested benchmarks. Our method effectively bridges the gap between deep representation learning and low-latency application requirements for CTDGs.
@@ 1 Introduction
Dynamic graphs are categorized into Discrete Time Dynamic Graphs (DTDGs) and Continuous Time Dynamic Graphs (CTDGs(Rossi et al., 2020). DTDGs are viewed as a sequence of snapshots at set intervals, while CTDGs are seen as a continuous stream of events, such as adding a new edge, which updates the graph’s structure with each occurrence. This paper aims to advance the state-of-the-art in representation learning for CTDGs.
Existing methods for handling CTDGs  (Dai et al., 2016; Kumar et al., 2019; Xu et al., 2020; Rossi et al., 2020) often face computational and memory constraints, leading to high-latency inference, thus limiting their practicality for real-time applications. Approaches such as asynchronous operation and truncated backpropagation have been employed to mitigate these issues, but they introduce compromises in representation accuracy and the learning of long-term dependencies (Rossi et al., 2020; Wang et al., 2021).
This paper introduces a novel architecture for the representation learning of CTDGs, designed to overcome existing limitations and provide low-latency, efficient representation learning. Our approach employs forward-mode automatic differentiation, specifically real-time recurrent learning (RTRL(Williams & Zipser, 1989), within a customized recurrent cell structure. This enables low-latency inference, efficient computation, and optimized memory usage, while preserving representation accuracy and the ability to capture long-term dependencies.
The contributions of this work are as follows:
@@
  • We benchmark DGS against state-of-the-art methods, in node classification and link prediction tasks, demonstrating on par predictive performance and up to 12x faster inference on our benchmarks, see Sections 4.
@@ 2 Background: Overview of Automatic Differentiation Modes
Temporal models, such as recurrent neural networks (RNNs) and GNNs for temporal graphs, pose specific challenges for AD due to their memory-intensive requirements in backpropagation. The memory complexity for storing intermediate states across timestamps significantly impacts computational efficiency. For instance, in an RNN with sequence length $l$ and state size $d$, reverse-mode AD exhibits computational and memory complexities of $\mathcal{O}(l\times d^{2})$, posing scalability issues for long sequences (Baydin et al., 2018). To mitigate these challenges, truncated backpropagation through time (TBPTT) optimizes resource usage by limiting the backpropagation horizon, thus reducing both computational and memory demands. However, TBPTT’s constraint on the temporal horizon restricts its ability to capture long-term dependencies, impacting model performance over extended sequences (Williams & Peng, 1990).
Forward-mode AD, exemplified in real-rime recurrent learning (RTRL), offers an alternative by facilitating online updates of the parameters, which is particularly advantageous for models requiring the retention of information over extended durations (or sequence length). Despite its benefits for capturing long-term dependencies with reduced memory overhead ($\mathcal{O}(d^{2})$), RTRL’s computational demand ($\mathcal{O}(l\times d^{4})$) limits its practicality in large-scale networks (Williams & Zipser, 1989; Cooijmans & Martens, 2019).
In summary, while reverse mode AD is challenged by memory intensity, TBPTT presents a compromise by reducing memory and computational needs at the expense of long-term dependency capture. Forward mode AD addresses both long-term retention and memory efficiency but is restricted by its computational complexity.
@@ 3 Method
In this section we detail our low latency node representation learning method, namely Deep-Graph-Sprints (DGS). We start by explaining the main components that form its architecture and then we detail each one of them. Furthermore, we detail the training paradigm that distinguishes our method, highlighting its enhanced memory efficiency and ability to capture long-term dependencies compared to existing approaches. Additionally, we explain our method during inference, demonstrating how it achieves low latencies.
@@ 3.1.1 Embedding Recurrent Component (ER):
$\vec{S}_{t}=\beta\vec{S}_{t-1}+(1-\beta)\left((1-\alpha)\vec{\delta}(F_{t})+\alpha\vec{S}^{*}_{t-1}\right)$ (1)
Here, $\vec{S}_{t}$ denotes the state of a node at time $t$, evolving from its previous state $\vec{S}_{t-1}$ and the state of its immediate neighbor $\vec{S}^{*}_{t-1}$, while also incorporating the current edge’s features $F_{t}$, encoded in histograms through the $\vec{\delta}$ function. The coefficients $\alpha$ and $\beta$ are scalar forgetting factors that modulate the impact of neighborhood and past information on the current state.
Although Graph-Sprints demonstrates rapid processing capabilities and performs on par with leading techniques, it faces several limitations in practical applications. These include the complex and time-consuming tuning processes required for its feature extraction and decision-making components. For instance, to tune the forgetting coefficients or to define the histogram bin edges of the $\vec{\delta}$ function for every feature. Moreover, the model’s expressivity is constrained by the uniform application of scalar forgetting coefficients across all features, limiting its ability to capture the unique temporal dynamics of each feature. Finally, the use of large number of histogram bins significantly increases the memory requirements especially in datasets with many features. In contrast, our proposed methodology, while drawing inspiration from Graph-Sprints, goes beyond traditional feature engineering by employing a dynamic learning mechanism for embeddings. The state update equation in our methodology is illustrated in Equation 2:
$\vec{S}_{t}=\vec{\beta}\odot\vec{S}_{t-1}+(1-\vec{\beta})\odot\left((1-\vec{\alpha})\odot\left(\operatorname*{\mathchoice{\Big{\|}}{\big{\|}}{\|}{\|}}_{i=1}^{m}\vec{\sigma}(W_{i}F_{t})\right)+\vec{\alpha}\odot\vec{S}^{*}_{t-1}\right)$ (2)
@@
Inspired by the Graph-Sprints histogram-based feature representations, we utilize the softmax function ($\vec{\sigma}$) to achieve analogous representations.Moreover, we employ softmax temperature scaling (Guo et al., 2017) where higher temperatures result in softer distributions. DGS enhances model expressiveness and optimizes memory usage by incorporating multiple softmax functions, each applied to a segment of the product between the embedding matrix $W$ and the feature vector $F_{t}$. The notation $\operatorname*{\mathchoice{\Big{\|}}{\big{\|}}{\|}{\|}}_{i=1}^{m}$ denotes the concatenation of the results obtained by applying the $m$ softmax functions. Each function is applied to the product of the $i$-th portion of the embedding matrix, $W_{i}$, and the features values $F_{t}$, as illustrated in Figure 2. This strategy not only aids in reducing computational and memory demands by lowering the Jacobian’s dimensionality but also introduces a modular structure akin to multi-head attention mechanisms in transformers (Vaswani et al., 2017), this technique has the potential to learn different information per softmax (further details in Section 3.2).
@@ 3.2 Training Process
In the NN component, number of learnable parameters varies based on model architecture, primarily involving the network’s weights. The parameters of the NN component are optimized using backpropagation, and to implement that we also leverage the functionalities of PyTorch.
As a result, the architecture of the DGS method employs a mixed-mode AD approach, as illustrated in Figure 3.
@@ 4.1 Experimental Setup
We compare the performance of our method against several baselines. The first simple baseline, called Raw, trains a machine learning model using only raw edge features. Another baseline is Graph-Sprints (Eddin et al., 2023) a graph feature engineering method, which we refer to by GS. The GS baseline uses the same ML classifier used by the Raw baseline but diverges in the features used for training— GS employs Graph-Sprints histograms, whereas Raw employs the raw edge features.
An additional baseline set comprises state-of-the-art GNN methods, specifically TGN (Rossi et al., 2020). Our TGN implementation, based on the default PyTorch Geometric implementation, differs from the original paper by restricting within-batch neighbor sampling, for a more realistic scenario.
@@
We leverage five different datasets, all CTDGs and labeled. Each dataset is split into train, validation, and test sets respecting time (i.e., all events in the train are older than the events in validation, and all events in validation are older than the events in the test set). Three of these datasets are publicly available (Kumar et al., 2019) from the social and education domains. In these three datasets, we adopt the identical data partitioning strategy employed by the baseline methods we compare against, which also utilized these datasets.
The other two datasets are real-world banking datasets from the AML domain. Due to privacy concerns, we can not disclose the identity of the FIs nor provide exact details regarding the node features. We refer to the datasets as FI-A and FI-B. The graphs in this use case are constructed by considering the accounts as nodes and the money transfers between accounts as edges. Table 1 shows the details of all the used datasets.
Wikipedia Mooc Reddit FI-A FI-B
#Nodes 9,227 7,047 10,984 $\approx$400,000 $\approx$10,000
#Edges 157,474 411,749 672,447 $\approx$500,000 $\approx$2,000,000
Positive labels 0.14% 0.98% 0.05% 2-5% 20-40%
Duration 30 days 29 days 30 days $\approx$300 days $\approx$600 days
Used split (%) 75-15-15 60-20-20 75-15-15 60-10-30 60-10-30
@@ 4.4 Inference Runtime
DGS has a primary goal of achieving reduced inference times. Comparative latency assessments were conducted amongst DGS, Graph-Sprints, and baseline GNN models. These assessments involved processing 200 batches, each containing 200 events, across distinct datasets (Wikipedia, Mooc, and Reddit) for node classification task. The average times were computed over 10 iterations. Tests were performed on a Linux PC equipped with 24 Intel Xeon CPU cores (3.70GHz) and an NVIDIA GeForce RTX 2080 Ti GPU (11GB).
As depicted in Figure 4, DGS exhibited significant speed advantages. On the Reddit dataset, it was more than ten times faster than the TGN-attn GNN baseline. For the smaller datasets, this speed enhancement ranged approximately between 5 and 6 times, while maintaining a competitive speed with the low latency GS baseline. Notably, the runtime of DGS remained stable and was not influenced by the number of edges in the graph, as demonstrated in Figure 4. In the Wikipedia and Reddit datasets, DGS consistently took 0.24 seconds. In contrast, TGN-attn exhibited a runtime increase from 1.2 seconds in the Wikipedia dataset to approximately 3 seconds in the Reddit dataset. This stability suggests that we may obtain higher speed gains in larger graphs, especially considering that inference times in the TGN-attn baseline are impacted by the number of graph edges. Moreover, When benchmarked against other GNNs baselines (TGN-ID, Jodie), DGS consistently demonstrated significantly lower inference latency.
@@ 4.5 Ablation Study
The initial variant, designated as DGS-s, represents the most basic approach wherein scalar parameters $\alpha$ and $\beta$ are learned. Moreover, instead of employing a learnable embedding matrix, DGS-s adopts the static embedding function utilized by Graph-Sprints. The subsequent variant, DGS-v, retains the fixed embedding function but transitions to vectorized parameters, $\vec{\alpha}$ and $\vec{\beta}$. This modification is intended to explore the effects of increasing these parameters’ complexity on the performance of the model. The final variant, the principal model described herein as DGS, not only incorporates vectorized parameters $\vec{\alpha}$ and $\vec{\beta}$ but also integrates a learnable embedding matrix. This approach aims to assess the impact of learnable feature embedding matrix.
@@ 5 Related Work
Graph representation learning is essential for converting complex graph structures into embeddings usable by machine learning models. This section provides an overview of the existing algorithms. Existing methods can be categorized into random-walk based, aggregation based, and other approaches.
5.1 Random-walk Based Techniques
Random-walk based techniques utilize random walks to approximate the inherent structure of the network. Nodes exhibiting similar random walks often indicate analogous structural roles within the network. Traditional random-walk methods like, DeepWalk (Perozzi et al., 2014), LINE (Tang et al., 2015), Node2vec (Grover & Leskovec, 2016), Struc2vec (Ribeiro et al., 2017), Metapath2vec (Dong et al., 2017), Role2vec (Ahmed et al., 2019), learn representation of static graphs. Adaptations for DTDG include (Sajjad et al., 2019), while Node2bits (Jin et al., 2019) integrates temporal information by segmenting random walks into distinct time windows. Further developments, such as CTDNE (Lee et al., 2020; Nguyen et al., 2018), were introduced to provide time-aware embeddings, augmenting the Node2vec (Grover & Leskovec, 2016) paradigm for CTDGs. These methodologies treat the graph as a stream of edges and perform temporal walks from seed nodes chosen through a temporally-biased distribution.
A major limitation of random-walk-based approaches is their computational and memory complexities as highlighted by Xia et al. (2019), there exists techniques to aleviate this like B_LIN (Tong et al., 2006), METIS (Karypis & Kumar, 1997)RWDISK (Sarkar & Moore, 2010), and Graph-Sprints (Eddin et al., 2023), which is a graph feature engineering approach that provide an approximation of random walks for low latency graph feature extraction. For a comprehensive review of random-walk-based applications and the challenges they pose, the reader is directed to (Xia et al., 2019).
5.2 K-hop Neighborhood Based Methods
More recently, we see a surge in deep learning algorithms for graph representation learning. Several approaches leverage GNNs to learn functions that generate node embeddings (Hamilton et al., 2017a; Ying et al., 2018; Yang et al., 2020; Zhu et al., 2020).
Most GNN-based methods require a $k$-hop neighborhood on which message-passing operations lead to node embeddings. To deal with CTDGs, a simple approach is to consider a series of discrete snapshots of the graph over time, on which static methods are applied. Such approaches however do not take time properly into account and several works propose techniques to alleviate this issue (Sankar et al., 2020; Goyal et al., 2018; Jin et al., 2022). To better deal with CTDGs, other works focus on including time-aware features or inductive biases into the architecture. DeepCoevolve (Dai et al., 2016) and Jodie (Kumar et al., 2019) train two RNNs for bipartite graphs, one for each node type. Importantly, the previous hidden state of one RNN is also added as an input to the other RNN. In this way, the two RNNs interact, in essence performing single-hop graph aggregations. TGAT (Xu et al., 2020) proposes to include temporal information in the form of time encodings, while TGN (Rossi et al., 2020) extends this framework and also includes a memory module taking the form of an RNN. Jin et al. (2020) replace the discrete-time recurrent network of TGN with a NeuralODE modeling the continuous dynamics of node embeddings.
GNNs are instrumental in analyzing vast time series data, but adapting them to large datasets is challenging due to memory constraints. While various sampling strategies exist, merging them with temporal data remains complex. Enhancing GNN’s scalability for real-time applications is a critical research area (Jin et al., 2023). For a comprehensive review on GNN-based approaches for time series analysis we refer the reader to the survey by Jin et al. (2023).
5.3 Low-latency methods

### v3

@@ Abstract
Continuous-time dynamic graphs (CTDGs) are essential for modeling interconnected, evolving systems. Traditional methods for extracting knowledge from these graphs often depend on feature engineering or deep learning. Feature engineering is limited by the manual and time-intensive nature of crafting features, while deep learning approaches suffer from high inference latency, making them impractical for real-time applications. This paper introduces Deep-Graph-Sprints (DGS), a novel deep learning architecture designed for efficient representation learning on CTDGs with low-latency inference requirements. We benchmark DGS against state-of-the-art (SOTA) feature engineering and graph neural network methods using five diverse datasets. The results indicate that DGS achieves competitive performance while inference speed improves between 4x and 12x compared to other deep learning approaches on our benchmark datasets. Our method effectively bridges the gap between deep representation learning and low-latency application requirements for CTDGs.
@@ 1 Introduction
Dynamic graphs are categorized into Discrete Time Dynamic Graphs (DTDGs) and Continuous Time Dynamic Graphs (CTDGs(Rossi et al., 2020). DTDGs are viewed as a sequence of snapshots at set intervals, while CTDGs are seen as a continuous stream of events, such as adding a new edge, which updates the graph’s structure with each occurrence. This paper aims to advance the state-of-the-art (SOTA) in representation learning for CTDGs.
Existing methods for handling CTDGs  (Dai et al., 2016; Kumar et al., 2019; Xu et al., 2020; Rossi et al., 2020) often face computational constraints, leading to high-latency inference, thus limiting their practicality for real-time applications. Approaches such as asynchronous operation and truncated backpropagation have been employed to mitigate these issues, but they introduce compromises in representation accuracy and the learning of long-term dependencies (Rossi et al., 2020; Wang et al., 2021).
This paper introduces a novel architecture for the representation learning of CTDGs, designed to overcome existing limitations and provide low-latency, efficient representation learning. Our approach employs forward-mode automatic differentiation, specifically real-time recurrent learning (RTRL(Williams & Zipser, 1989), within a customized recurrent cell structure. This enables low-latency inference, efficient computation, and optimized memory usage, while preserving representation accuracy and the ability to capture long-term dependencies. The contributions of this work are as follows:
@@
  • We benchmark DGS against SOTA methods, in node classification and link prediction tasks, demonstrating on par predictive performance while achieving significantly faster inference speed—ups up to 12x faster than TGN-attn (Rossi et al., 2020), and up to 8x faster than both TGN-ID (Rossi et al., 2020) and Jodie (Kumar et al., 2019). Detailed results are provided in Section 4.
@@ 2 Background: Overview of Automatic Differentiation Modes
Temporal models, such as recurrent neural networks (RNNs) and GNNs for temporal graphs, pose specific challenges for backpropagation due to their memory-intensive requirements. The memory complexity for storing intermediate states across a history significantly impacts the feasibility of full backpropagation. For instance, in an RNN with sequence length $l$ and state size $d$, backpropagation-through-time exhibits computational and memory complexities of $\mathcal{O}(l\times d^{2})$, posing scalability issues for long sequences (Baydin et al., 2018). To mitigate these challenges, truncated backpropagation through time (TBPTT) optimizes resource usage by limiting the backpropagation horizon, thus reducing both computational and memory demands. However, TBPTT’s constraint on the temporal horizon restricts its ability to capture long-term dependencies, impacting model performance over extended sequences (Williams & Peng, 1990).
Forward-mode AD, exemplified in real-time recurrent learning (RTRL), offers an alternative by facilitating online updates of the parameters, which is particularly advantageous for models requiring the retention of information over extended durations (or sequence length). Despite its benefits for capturing long-term dependencies with reduced memory overhead ($\mathcal{O}(d^{2})$), RTRL’s computational demand ($\mathcal{O}(d^{4})$) limits its practicality in large-scale networks (Williams & Zipser, 1989; Cooijmans & Martens, 2019).
In summary, while full backpropagation through time has high memory demands, TBPTT presents a compromise by reducing memory and computational needs at the expense of long-term dependency capture. Forward mode AD (RTRL in our case) addresses both long-term retention and memory efficiency but is restricted by its computational complexity.
To address these limitations, our approach leverages RTRL, thus benefiting from its low memory footprint, and combines it with a custom architecture that reduces its computational complexity to match that of backpropagation. This design effectively mitigates the computational challenges typically associated with forward-mode AD, and allows the model to capture long-term dependencies unlike TBPTT.
@@ 3 Method
In this section we detail our low latency node representation learning method, namely Deep-Graph-Sprints (DGS). We start by explaining the main components that form its architecture and then we detail each one of them. Furthermore, we detail the training paradigm that distinguishes our method, highlighting its memory demands and ability to capture long-term dependencies compared to existing approaches through its RTRL-based approach. Additionally, we explain our method during inference, demonstrating how it achieves low latencies.
@@ 3.1.1 Embedding Recurrent Component (ER):
$\vec{S}_{t}=\beta\vec{S}_{t-1}+(1-\beta)\left((1-\alpha)\vec{\delta}(\vec{F_{t}})+\alpha\vec{S}^{*}_{t-1}\right)$ (1)
Here, $\vec{S}_{t}$ denotes the state of a node at time $t$, evolving from its previous state $\vec{S}_{t-1}$ and the state of its immediate neighbor $\vec{S}^{*}_{t-1}$, while also incorporating the current edge’s features $\vec{F_{t}}$, encoded through the $\vec{\delta}$ function, which bucketizes these features. The $\vec{\delta}$ function maps numerical features into a one-hot encoded vector based on predetermined buckets (e.g., buckets corresponding to specific percentiles). This bucketization provides a deterministic representation of the input features, which requires manual tuning to define the buckets and results in sparse, high-dimensional feature vectors. The coefficients $\alpha$ and $\beta$ are scalar forgetting factors that modulate the impact of neighborhood and past information on the current state.
Although Graph-Sprints demonstrates rapid processing capabilities and performs on par with leading techniques, it faces several limitations in practical applications. These include the complex and time-consuming tuning processes required for its feature extraction and decision-making components. For instance, to tune the forgetting coefficients or to define the bucketed features edges of the $\vec{\delta}$ function for every feature. Moreover, the model’s expressivity is constrained by the uniform application of scalar forgetting coefficients across all features, limiting its ability to capture the unique temporal dynamics of each feature. Finally, the use of large number of buckets per feature significantly increases the memory requirements especially in datasets with many features. In contrast, our proposed methodology, while drawing inspiration from Graph-Sprints, goes beyond traditional feature engineering by employing a dynamic learning mechanism for embeddings. The state update equation in our methodology is illustrated in Equation 2:
$\vec{S}_{t}=\vec{\beta}\odot\vec{S}_{t-1}+(1-\vec{\beta})\odot\left((1-\vec{\alpha})\odot\left(\operatorname*{\mathchoice{\Big{\|}}{\big{\|}}{\|}{\|}}_{i=1}^{m}\vec{\sigma}(W_{i}\vec{F_{t}})\right)+\vec{\alpha}\odot\vec{S}^{*}_{t-1}\right)$ (2)
@@
Inspired by the Graph-Sprints feature representations, where the buckets that belong to the same feature sum to one, we utilize the softmax function ($\vec{\sigma}$) to achieve analogous representations.Moreover, we employ softmax temperature scaling (Guo et al., 2017) where higher temperatures result in softer distributions. DGS enhances model expressiveness and optimizes memory usage by incorporating multiple softmax functions, each applied to a segment of the product between the embedding matrix $W$ and the feature vector $\vec{F_{t}}$. The notation $\operatorname*{\mathchoice{\Big{\|}}{\big{\|}}{\|}{\|}}_{i=1}^{m}$ denotes the concatenation of the results obtained by applying the $m$ softmax functions. Each function is applied to the product of the $i$-th portion of the embedding matrix, $W_{i}$, and the features values $\vec{F_{t}}$, as illustrated in Figure 2. This strategy not only aids in reducing computational and memory demands by limiting the dependency of each state element to a specific segment of the embedding matrix and thereby lowering the Jacobian’s dimensionality but also introduces a modular structure.
@@ 3.1.2 Neural Network (NN)
Although Equation 1 (GS) and Equation 2 (DGS) present similarities. There are several aspects that distinguish Graph-Sprints and DGS methods. GS is a feature engineering approach with fixed embeddings, while DGS employs deep learning with learnable embeddings. Moreover, parameter optimization in GS is separate, whereas DGS allows end-to-end optimization. Table 4 compares both methods.
@@ 3.2 Training Process
In the NN component, number of learnable parameters varies based on model architecture, primarily involving the network’s weights. The parameters of the NN component are optimized using backpropagation, and to implement that we also leverage the functionalities of PyTorch. As a result, the architecture of the DGS method employs a mixed-mode AD approach, as illustrated in Figure 3.
@@
Compared to the GS approach, allowing the key parameters ($\vec{\alpha}$, $\vec{\beta}$, and $W$) to be learned from the data overcomes the limitations of separate tuning processes, thereby simplifying the training procedure. Additionally, this enables the use of vectorized forgetting coefficients instead of scalar ones, which significantly enhances the model’s expressivity.
@@ 3.3 Inference Process
It is important to note that the DGS method is fully online and supports both single-sample and mini-batch inference, offering flexibility depending on the scenario. Mini-batch inference, when suitable, improves both training and inference speeds even further. In our experiments (Section LABEL:sec:deep_graph_sprints_results), we utilize mini-batch inference to ensure comparability with other experimental setups used in state-of-the-art studies.
@@ 4.1 Experimental Setup
We compare the performance of our method against several baselines. The first simple baseline, called Raw, trains a machine learning model using only raw edge features. Another baseline is Graph-Sprints (Eddin et al., 2023) a graph feature engineering method, which we refer to by GS. The GS baseline uses the same ML classifier used by the Raw baseline but diverges in the features used for training— GS employs Graph-Sprints encoded features, whereas Raw employs the raw edge features.
An additional baseline set comprises SOTA GNN methods, specifically TGN (Rossi et al., 2020). Our TGN implementation, based on the default PyTorch Geometric implementation, differs from the original paper by restricting within-batch neighbor sampling, for a more realistic scenario.
@@
We leverage five different datasets, all CTDGs and labeled. Each dataset is split into train, validation, and test sets respecting time (i.e., all events in the train are older than the events in validation, and all events in validation are older than the events in the test set). Three of these datasets are public (Kumar et al., 2019) from the social and education domains. In these three datasets, we adopt the identical data partitioning strategy employed by the baseline methods we compare against, which also utilized these datasets. The other two datasets are real-world banking datasets from the AML domain. Due to privacy concerns, we can not disclose the identity of the FIs nor provide exact details regarding the node features. We refer to the datasets as FI-A and FI-B. The graphs in this use case are constructed by considering the accounts as nodes and the money transfers between accounts as edges. Table 5 shows the details of all the used datasets.
@@ 4.2 Node classification
In the node classification task, given the dataset characteristics detailed in Table 5, we address binary classification with class imbalance. To evaluate performance, we calculate the area under the ROC curve (ROC-AUC), referred to as AUC for brevity, by plotting the True Positive Rate (TPR) against the False Positive Rate (FPR) across various classification thresholds, and then computing the area under the curve. The AUC is calculated using the ’sklearn’ library in python.
@@ 4.4 Inference Runtime
DGS has a primary goal of achieving reduced inference times. Comparative latency assessments were conducted amongst DGS, Graph-Sprints, and baseline GNN models. These assessments involved processing 200 batches, each containing 200 events, across distinct datasets (Wikipedia, Mooc, and Reddit) for node classification task. The average times were computed over 10 iterations. Tests were performed on a Linux PC equipped with 24 Intel Xeon CPU cores (3.70GHz) and an NVIDIA GeForce RTX 2080 Ti GPU (11GB). Note that all experiments, including those for link prediction and node classification mentioned in the previous sections, used the same machine.
As depicted in Figure 4, DGS exhibited significant speed advantages. On the Reddit dataset, it was more than ten times faster than the TGN-attn GNN baseline. For the smaller datasets, this speed enhancement ranged approximately between 5 and 6 times, while maintaining a competitive speed with the low latency GS baseline. Notably, the runtime of DGS remained stable and was not influenced by the number of edges in the graph, as demonstrated in Figure 4. In the Wikipedia and Reddit datasets, DGS consistently took 0.24 seconds. In contrast, TGN-attn exhibited a runtime increase from 1.2 seconds in the Wikipedia dataset to approximately 3 seconds in the Reddit dataset. This stability suggests that we may obtain higher speed gains in larger or denser graphs, especially considering that inference times in the TGN-attn baseline are impacted by the number of graph edges. Moreover, When benchmarked against other GNNs baselines (TGN-ID, Jodie), DGS consistently demonstrated significantly lower inference latency.
@@ 4.5 Ablation Study
The initial variant, designated as DGS-s, represents the most basic approach wherein scalar parameters $\alpha$ and $\beta$ are learned. Moreover, instead of employing a learnable embedding matrix, DGS-s adopts the static embedding function utilized by Graph-Sprints. The subsequent variant, DGS-v, retains the fixed embedding function but transitions to vectorized parameters, $\vec{\alpha}$ and $\vec{\beta}$. This modification aims to explore the effects of increasing these parameters’ complexity on the performance of the model. The final variant, referred to as DGS, not only incorporates vectorized parameters $\vec{\alpha}$ and $\vec{\beta}$ but also integrates a learnable embedding matrix. This approach aims to assess the impact of learnable feature embedding matrix.
@@
Further ablation studies demonstrating the advantages of using forward-mode AD and softmax normalization are detailed in Appendix A.4. Moreover, the differences in inference speed are discussed in Appendix A.5
@@ 5 Related Work
Graph representation learning is essential for converting complex graph structures into embeddings usable by machine learning models. This section provides an overview of the existing algorithms. Additionally, given the focus of this paper, we explore approaches for low latency in graph representation learning.
5.1 Graph Representation Learning
Most existing graph representation learning methods focus on static graphs, thereby neglecting temporal dynamics (Perozzi et al., 2014; Tang et al., 2015; Grover & Leskovec, 2016; Hamilton et al., 2017a; Ying et al., 2018). Dynamic graphs, which evolve over time, introduce additional complexities. A common approach is to use DTDGs by considering the dynamic graph a series of discrete snapshots and apply static methods (Sajjad et al., 2019), but this approach fails to capture the full spectrum of temporal dynamics.
To address this limitation, more advanced techniques have been developed to better handle CTDGs. These methods include incorporating time-aware features or inductive biases into the architecture(e.g.,  (Nguyen et al., 2018; Jin et al., 2019; Lee et al., 2020; Rossi et al., 2020)).
For instance, methods like DeepCoevolve(Dai et al., 2016) and Jodie(Kumar et al., 2019) train two recurrent neural networks (RNNs) for bipartite graphs, one for each node type. In these models, the previous hidden state of one RNN is also used as an input to the other RNN, allowing interaction between the two and effectively performing single-hop graph aggregations.
TGAT(Xu et al., 2020) introduces temporal information through time encodings, enhancing the model’s ability to capture dynamic changes. TGN(Rossi et al., 2020) extends this approach by incorporating a memory module in the form of an RNN, providing a more robust framework for handling temporal data. Further refinement is seen in Jin et al. (2020), where the discrete-time recurrent network of TGN is replaced with a NeuralODE, modeling the continuous dynamics of node embeddings for more accurate representations.
The methods described above either leverage random-walks or graph neural networks (GNNs) to extract neighborhood information and understand graph structure. Random-walk-based methods are often hindered by high computational and memory costs, as noted by Xia et al. (2019). Solutions to mitigate these challenges include techniques such as B_LIN(Tong et al., 2006), METIS(Karypis & Kumar, 1997), and RWDISK (Sarkar & Moore, 2010), which offer approximations of random walks.
GNNs are powerful for representation learning of graphs, but adapting them to extensive datasets poses challenges. When the graph does not fit in memory, sampling the neighbors of a node may necessitate costly disk reads. While various sampling strategies have been proposed, integrating these with temporal data is complex. Enhancing the scalability of GNNs for real-time applications remains a critical area of ongoing research (Jin et al., 2023).
5.2 Low-latency Graph Representation Learning